step by step mechanism of fischer esterification
O=CO S-OH ROH OR + H₂O OH microwave
Answer #1

= S-OH O= + ROH i Ho + ROH -OH OR + H₂O Micro- - wave Mechanism: 0 I H-O O=y= höh 10: -O.H 8.no RÖH I ö-H H-o: :00 0 TOH R 0-